10.6 A company claims that its batteries have a mean life of 100 hours. You try to verify this for a sample of size 21 with mean 97 hours and variance 9 hours.


1
Expert's answer
2021-07-13T05:21:20-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=100"

"H_1:\mu\not=100"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05," "df=n-1=21-1=20" degrees of freedom and the critical value for a two-tailed test is "t_c=2.086."

The rejection region for this two-tailed test is "R=\{t: |t|>2.086\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{97-100}{9/\sqrt{21}}=-1.5275"

Since it is observed that "|t|=1.5275<2.086=t_c," it is then concluded that the null hypothesis is not rejected.


Using the P-value approach: The p-value for "\alpha=0.05, df=20" degrees of freedom, two-tailed, "t=-1.5275" is "p=0.142298," and since "p=0.142298>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than "100," at the "\alpha=0.05" significance level.